# Exploring Holomorphic Retracts

### v1

@@ 2. Preliminaries
We recall certain basic and general facts about retracts which would needless to say, apply to our specific case and put it in good perspective, as well. Let us first recall that a self-map $F$ of a set $X$ is called a retraction if $F\circ F=F$; note that this idempotent condition on $F$ means that $F$ acts as the identity map on its range $Z=F(X)$. So, for any Hausdorff topological space $X$, its retracts, which are the images of continuous retractions, are closed subsets of $X$. If $X$ is connected, then so is any of its retracts $Z$. When $X$ is furthermore contractible i.e., the space $X$ deformation retracts to a point then by composing any such deformation retraction with the retraction map $F$, gives a homotopy within $Z$ that reduces $Z=F(X)$ to a point, meaning that if $X$ is contractible then so are all of its retracts. If $X$ even happens to be a convex domain in some real topological vector space, then the retracts of $X$ are also strong deformation retracts. We are ofcourse interested in holomorphic retracts: if $X$ is a complex manifold and $F$ is a holomorphic self-map of $X$ which is idempotent (i.e., $F\circ F=F$) then, we say that $F$ is a holomorphic retraction and its image $Z=F(X)$ a holomorphic retract; as we shall only be talking about holomorphic retracts, we shall most often drop the adjective ‘holomorphic’. Moreover as mentioned in the introduction, our focus will be to determine the form of the retracts $Z=F(X)$ rather than the retractions $F$. A first basic and general fact about holomorphic retracts that is helpful to have in the background for the discussion in the sequel is laid down in the following well-known theorem which says that they are not just analytic varieties but actually (connected) manifolds i.e., are smooth and have the same dimension throughout.
@@
Continuing with the proof theorem 2.1, we shall now show that in a neighbourhood of any given $p\in Z$, $Z$ is the level set of a holomorphic map whose derivative has constant rank near $p$; the implicit mapping theorem will then render the claim of the theorem. To this end then, we fix an arbitrary $p\in Z$ and passing to a holomorphic chart of the complex manifold $X$ near $p$, we may as well assume (as already mentioned), that we are in the complex Euclidean space $\mathbb{C}^{N}$, which allows us to define $\beta(x)=x-\rho(x)$ for $x$ near $p$, whose zero set is precisely $Z$ (all in the charts near $p$, to be precise). To show that the constancy of the rank of the derivative $D\beta(x)$ as mentioned, we first consider the derivative of $\rho$. The idempotence of the linear operator $D\rho(p)$ given by the above lemma, guarantees its diagonalizability as well. So, after a linear change of coordinates, $D\rho(p)$ is a diagonal matrix with diagonal entries either $1$ or 00. Let $r$ be the number of $1$’s i.e., $r={\rm rank}(D\rho(p))$. We may as well assume by a permutation of coordinates if needed, that the first principal $r\times r$ submatrix of $D\rho(p)$ is the identity matrix of order $r$, which by continuity of the determinant function implies that the determinant of the same principal minor of the nearby derivative matrices $D\rho(x)$ for all $x$ in some neighbourhood (contained in the chart) about $p$, is non-zero as well. This just means the well-known lower semi-continuity of the rank function, which when applied to $x\mapsto{\rm rank}(D\rho(x))$ implies that the rank of $D\rho(x)$ is atleast $r$ for all $x$ in some neighbourhood of $p$. Now, observe that the same arguments may be applied to $D\beta(x)$ as well. Firstly, note that $D\beta(p)=I-D\rho(p)$ where $I$ is the identity matrix of size $n$, is also an idempotent matrix, indeed diagonal with diagonal entries only $1$ or 00 and with the number of zeros being exactly $n-r$. In particular, ${\rm rank}(D\beta(p))=n-r$ and the lower semicontinuity of the rank applied to $x\mapsto{\rm rank}(D\beta(x))$ implies that the rank of $D\beta(x)$ is atleast $n-r$ in a neighbourhood of $p$. Now, we have a similar lower bound on the rank of $D\rho(x)$ as already noted, which is tempting to be translated this into an bound in-terms of $\beta$ instead of $\rho$; specifically, the upper bound ${\rm rank}(D\beta(x))\leq n-r$ in the some neighbourhood of $p$. While this is correct for $x=p$ and even along $Z$, it is not at all apriori clear as to why this bound continued to hold in an open neighbourhood of $p$. The reason for detailing this flawed argument is that it is then easy to finish the proof, within standard elementary ideas of calculus on differential manifolds. Indeed firstly, the foregoing pair of bounds when put together, only means that the rank of $D\beta(x)$ is constant in some neighbourhood of $p$, leading to the conclusion by the constant rank theorem that $Z$ is a complex manifold of dimension $r$ throughout, as we already know $Z=\rho(X)$ is connected but alas, this part of the argument is flawed, as pinpointed. This finishes the discussion of the proof of 2.1. Let us round this by noting a couple of very easy but noteworthy facts. Retracts of compact complex manifolds are compact complex submanifolds. Retracts of Stein manifolds are Stein and likewise, the property of a complex manifold being an Oka manifold, is preserved under passage to a retract thereof, as well.
Another general fact to keep in the background is the following; we include its proof, despite its obviousness, owing to its use for the rest of this article.
@@
In passing, we may ofcourse note that by using the well-known theorem of Kaup–Upmeier [Kaup-Upm], we may replace the polydisc in the above by various other product domains which unlike polydiscs are neither convex nor homogeneous. Specifically, we may take a product $B_{1}\times\ldots\times B_{k}$ (with $k\geq 2$) of bounded balanced pseudoconvex domains with the property that one of the factors admits a non-linear holomorphic map into another, to ensure the existence of non-linear retracts, thereby the proof of the above corollary to go through. There are of-course better proofs and our reason for discussing this proof is to simultaneously make a few general remarks about non-linear retracts and about some simple ways of generating ample examples of retracts. We note that there are plenty of product domains of the kind just mentioned (satisfying the property about the existence of non-linear holomorphic maps, as mentioned above). To indicate a way of generating examples, let $E_{2m}:=\{(z,w)\;:\;|z|^{2m}+|w|^{2}<1\}$. For an arbitrary point $(z,w)\in E_{2m}$ note that $|z|<1$, thereby that, $|z^{2l}|+|w|^{2}<|z|^{2m}+|w|^{2}$ for every $l\geq m$. Consequently, $(z^{2l},w)$ lies in $E_{2l}$ for every $l\geq m$, implying that $(z,w)\mapsto(z^{2l},w)$, gives an example of a non-linear (in-fat proper) holomorphic mapping as desired and more importantly thereby, a non-linear retract passing through the origin in the polyball $E_{2m}\times E_{2l}$ for every $l\geq m$. Needless to say, similar constructions can be done with $l_{p}$-balls and more; we shall not digress further along this line.
@@ 2.2. Notions of extremality and strict convexity
Let $K$ be the field $\mathbb{R}$ or $\mathbb{C}$. Let $\Omega\subset\mathbb{C}^{N}$ be an arbitrary domain and $a\in\partial{\Omega}$. We introduce a notion that captures the set of all directions along which one can move from $a$, at-least a little bit while staying within the boundary $\partial\Omega$. Towards this, define the face of $\partial\Omega$ at $a$ by
$F_{K}(a,\partial{\Omega})=\{u\in\mathbb{C}^{N}:\exists\epsilon>0,a+tu\in\partial{\Omega},\text{ for all }t\in K~{}with~{}|t|<\epsilon\}$
reduces to $\{0\}$. It is immediate that $F_{K}$ is closed under scalar multiplication. But a stronger claim that $F_{K}$ is a linear subspace of $\mathbb{C}^{N}$ requires more assumptions on $\Omega$. Indeed, if $\Omega$ is convex, then $F_{K}(a,\partial{\Omega})$ is $K$-linear subspace. A boundary point $a\in\partial{\Omega}$ is said to be a $K$-extremal boundary point for $\Omega$, if $F_{K}(a,\partial{\Omega})=0$; if the face reduces to $\{0\}$ for every $a\in\partial{\Omega}$, then we say that the domain $\Omega$ is $K$-extremal.
@@
$\vec{F}_{\mathbb{C}}(a,X)=\vec{F}_{\mathbb{R}}(a,X)\cap i\vec{F}_{\mathbb{R}}(a,X).$
It follows that if $\Omega$ is $\mathbb{R}$-extremal, then $\Omega$ is $\mathbb{C}$-extremal. But the converse is not true. To see this, consider $B_{\ell^{1}}:=\{(z,w)\in\mathbb{C}^{2}:|z|+|w|<1\}$; we shall presently proceed to showing that this is $\mathbb{C}$-extremal but not $\mathbb{R}$-extremal. Perhaps even more basically, we must remark that every domain in $\mathbb{C}$ is $\mathbb{C}$-extremal. But only strict convex domains are $\mathbb{R}$- extremal.
@@
We first show that $B_{\ell^{1}}$ is $\mathbb{C}$-extremal. To do this, let $a\in\partial B_{\ell^{1}}$. Suppose there exists $\epsilon_{0}>0,a\in\partial B_{\ell^{1}}$ and $u_{0}\in\mathbb{C}^{N}$ such that $a+tu_{0}\in\partial B_{\ell^{1}}$, for all $t\in\mathbb{C}$ with $|t|<\epsilon_{0}$.
@@
First we prove that the notion of rotundity and strictly convex are equivalent in a bounded balanced convex domain $\Omega$.
@@
Assume $\Omega$ is rotund. Let $a\in\partial\Omega$ and $(a+tu)\in\partial\Omega$ for all $t\in\mathbb{R}$ with $|t|<\epsilon$. Note that $\frac{1}{2}(a+(a+tu))=a+\frac{t}{2}u\in\partial\Omega$ for all $t\in\mathbb{R}$ with $|t|<\epsilon$. By proposition 5.1.2 of [Meggn], $a=a+tu$ for all $t\in\mathbb{R}$ with $|t|<\epsilon$ and this implies that $u=0$. Therefore, $\Omega$ is $\mathbb{R}$extremal.
Conversely assume that $\Omega$ is $\mathbb{R}$extremal. Let $z_{1},z_{2}\in\partial\Omega$.
@@
$\frac{z_{1}+z_{2}}{2}+t(z_{2}-z_{1})=(1/2-t)z_{1}+(1/2+t)z_{2}\in\partial{\Omega}~{}\text{for ~{}all}\ t\ \text{with}\ |t|<1/2$
Therefore $z_{2}-z_{1}\in F_{\mathbb{R}}(\frac{z_{1}+z_{2}}{2},\partial{\Omega})$. Since $\Omega$ is $\mathbb{R}$- extremal, $z_{2}=z_{1}$ and therefore $\Omega$ is rotund.
As the relation between the pair of notions gotten by taking $K=\mathbb{C}$ versus $K=\mathbb{R}$ will be basic importance in the sequel, we discuss this below. But one would perhaps like to know first the relation between strict convexity and extremality introduced above; as we do not find a quick reference for this, we include the details for this here.
@@
Suppose that $\Omega$ is strictly $\mathbb{R}$-convex. If possible, assume that $\Omega$ is not $\mathbb{R}$- extremal. Then by theorem 2.15, $\Omega$ is not rotund. By corollary 5.1.16 of [Meggn], there exists $\mathbb{R}$- linear functional $f$ supports more than one point. This implies that there exists $z_{0},w_{0}\in\mathbb{C}^{N}$ such that $f(z_{0})=f(w_{0})=\|f\|$. Take $\alpha:=\|f\|$. Define a $\mathbb{R}$- linear functional $g$ defined by $g(z)=\frac{1}{\alpha}f(z)$. Note that $\|g|\ =1$ and $g(z_{0})=g(w_{0})=1$. This contradicts the strict $\mathbb{R}$-convexity of $\Omega$. Therefore, $\Omega$ is $\mathbb{R}$- extremal.
Conversely assume that $\Omega$ is $\mathbb{R}$- extremal. Let $T$ be a $\mathbb{R}-$ linear map with $\|T\|=1$. Assume that there exists $z,w\in\overline{\Omega}$ such that $Tz=1=Tw$. Then $z,w\in\partial{\Omega}$, otherwise $T\left(\frac{z}{\|z\|}\right)=\frac{1}{\|z\|}>1$. Note that $T((1-t)z+tw)=1$. Then $(1-t)z+tw\in\partial{\Omega}$ for all $t\in[0,1]$. In particular, if $t=1/2$, we get $\frac{1}{2}(z+w)\in\partial{\Omega}$. Note then that
@@
Then $w-z\in F_{\mathbb{R}}(\frac{z+w}{2},\partial{\Omega})$. Since $\Omega$ is $\mathbb{R}$-extremal, then $w=z$. Therefore $\Omega$ is strictly $\mathbb{R}$-convex. ∎
@@
Conversely, assume that $\Omega$ is strictly $\mathbb{C}$-convex. We have to prove that $\Omega$ is strictly $\mathbb{R}$-convex. By above theorem, it suffices to show that $\Omega$ is $\mathbb{R}$-extremal. If possible assume that there exists $u\in\mathbb{C}^{N},~{}a\in\partial\Omega$ and an $\epsilon>0$ such that $a+tu\in\partial\Omega$ for all $t\in\mathbb{R}$ with $|t|<\epsilon$. Let $M=span_{\mathbb{R}}\{u\}$.

### v6

@@ 2. Preliminaries
We recall certain basic and general facts about retracts which would needless to say, apply to our specific cases and put it in good perspective, as well. While we would like to minimize repetition, we prefer to put the details on record to ensure a rigorous exposition. But here we repeat a few generalities from the introductory section for one last time, to cast this section systematically. Let us begin with the most general definition relevant to us: a self-map $\rho$ of a set $X$ is called a retraction if $\rho\circ\rho=\rho$; note that this idempotent condition on $\rho$ means that $\rho$ acts as the identity map on its range $Z=\rho(X)$. So, for any Hausdorff topological space $X$, its retracts, which are the images of continuous retractions, are closed subsets of $X$; henceforth Hausdorffness on $X$ will be in force. We regard $\rho$ as being non-trivial, if it is neither a constant map nor the identity. In addition to the contrasts between proper maps and retractions mentioned in the introduction, we note that non-trivial retraction mappings may very well fail to be closed maps notwithstanding their ranges being closed subsets always; indeed, as soon as $X$ is a non-compact complex manifold – the general setting of interest in this article – they always fail to be closed maps, despite the general fact that $\rho(X)$ is always closed as a subset of $X$, for every retraction map $\rho$ on $X$ as just mentioned (proofs detailed later below). However, the feature of being a quotient map is common to both retraction maps as well as proper maps. Getting back to the general case of a Hausdorff topological space $X$, if we add the assumption that $X$ is connected, then so are every of its retracts $Z$. When $X$ is furthermore contractible i.e., the space $X$ deformation retracts to a point then by composing any such deformation retraction with the retraction map $\rho$, gives a homotopy within $Z$ that reduces $Z=\rho(X)$ to a point, meaning that if $X$ is contractible then so are all of its retracts. If $X$ even happens to be a convex domain in some real topological vector space, then the retracts of $X$ are also strong deformation retracts; indeed, the homotopy $H(t,x)=(1-t)x+t\rho(x)$ substantiates this. We are ofcourse interested in holomorphic retracts: if $X$ is a complex manifold and $\rho$ is a holomorphic self-map of $X$ which is idempotent then, we say that $\rho$ is a holomorphic retraction and its image $Z=\rho(X)$ a holomorphic retract; as we shall only be talking about holomorphic retracts, we shall most often drop the adjective ‘holomorphic’. Moreover as mentioned in the introduction, our focus will be to determine the form of the retracts $Z=\rho(X)$ rather than the retractions $\rho$. A first basic and general fact about holomorphic retracts that is helpful to have in the background for the discussion in the sequel is laid down in the following well-known theorem which says that they are not just analytic varieties but actually (connected) manifolds i.e., are smooth and have the same dimension throughout.
@@
Continuing with the proof of theorem 2.1, we shall now show that in a neighbourhood of any given $p\in Z$, $Z$ is the level set of a holomorphic map whose derivative has constant rank near $p$; the implicit mapping theorem will then render the claim of the theorem. To this end then, we fix an arbitrary $p\in Z$ and passing to a holomorphic chart of the complex manifold $X$ near $p$, we may as well assume (as already mentioned), that we are in the complex Euclidean space $\mathbb{C}^{N}$, which allows us to define $\beta(x)=x-\rho(x)$ for $x$ near $p$, whose zero set is precisely $Z$ (all in the charts near $p$, to be precise). To show that the constancy of the rank of the derivative $D\beta(x)$ as mentioned, we first consider the derivative of $\rho$. The idempotence of the linear operator $D\rho(p)$ given by the above lemma, guarantees its diagonalizability as well. So, after a linear change of coordinates, $D\rho(p)$ is a diagonal matrix with diagonal entries either $1$ or 00. Let $r$ be the number of $1$’s i.e., $r={\rm rank}(D\rho(p))$. We may as well assume by a permutation of coordinates if needed, that the first principal $r\times r$ submatrix of $D\rho(p)$ is the identity matrix of order $r$, which by continuity of the determinant function implies that the determinant of the same principal minor of the nearby derivative matrices $D\rho(x)$ for all $x$ in some neighbourhood (contained in the chart) about $p$, is non-zero as well. This just means the well-known lower semi-continuity of the rank function, which when applied to $x\mapsto{\rm rank}(D\rho(x))$ implies that the rank of $D\rho(x)$ is atleast $r$ for all $x$ in some neighbourhood of $p$. Now, observe that the same arguments may be applied to $D\beta(x)$ as well. Firstly, note that $D\beta(p)=I-D\rho(p)$ where $I$ is the identity matrix of size $n$, is also an idempotent matrix, indeed diagonal with diagonal entries only $1$ or 00 and with the number of zeros being exactly $n-r$. In particular, ${\rm rank}(D\beta(p))=n-r$ and the lower semicontinuity of the rank applied to $x\mapsto{\rm rank}(D\beta(x))$ implies that the rank of $D\beta(x)$ is atleast $n-r$ in a neighbourhood of $p$. Now, we have a similar lower bound on the rank of $D\rho(x)$ as already noted, which is tempting to be translated this into an bound in-terms of $\beta$ instead of $\rho$; specifically, the upper bound ${\rm rank}(D\beta(x))\leq n-r$ in the some neighbourhood of $p$. While this is correct for $x=p$ and even along $Z$, it is not at all apriori clear as to why this bound continued to hold in an open neighbourhood of $p$. The reason for detailing this flawed argument is that it is then easy to finish the proof, within standard elementary ideas of calculus on differential manifolds. Indeed firstly, the foregoing pair of bounds when put together, only means that the rank of $D\beta(x)$ is constant in some neighbourhood of $p$, leading to the conclusion by the constant rank theorem that $Z$ is a complex manifold of dimension $r$ throughout, as we already know $Z=\rho(X)$ is connected but alas, this part of the argument is flawed, as pinpointed. This finishes the discussion of the proof of 2.1. Let us round this by noting a couple of very easy but noteworthy facts. Retracts of compact complex manifolds are compact complex submanifolds. Retracts of Stein manifolds are Stein and likewise, the property of a complex manifold being an Oka manifold, is preserved under passage to a retract thereof, as well.
Proposition 2.3.
Let $X$ be a non-compact complex manifold with $Z$ being any non-trivial retract of $X$. Then none of the retraction maps $\rho:X\to Z$ is a closed map.
Proof.
Choose any convergent sequence $(z_{n})$ in $Z$ of distinct points; as we already know $Z$ is closed, thereby the limit $z_{0}$ of the chosen sequence lies in $Z$. Our strategy will be to pick points $p_{n}$ in the fibers over $z_{n}$ such that the $p_{n}$’s form a discrete, thereby a closed subset $C$  of $X$. If we let $K$ denote the (infinite) compact set $K=\{z_{n}\}\cup\{z_{0}\}$ and note that $\rho(C)$ equals $K$ without its limit point, we would have proven the proposition. The possibility of choosing a sequence $p_{n}$ as desired above really requires the non-compactness in hypothesis. Indeed, consider the sequence of fibers of $\rho$ given by $F_{n}:=\rho^{-1}(z_{n})$. Note firstly that being non-trivial analytic sets, each of the $F_{n}$’s is non-compact and $\rho^{-1}(K\setminus\{z_{0}\})=\cup_{n=1}^{\infty}F_{n}$. We then just use the fact from general topology that: being second countable, locally compact and Hausdorff, $X$ admits an exhaustion by compact subsets $K_{n}$. In particular, since $X$ is not compact, the complements $K_{n}^{c}=X\setminus K_{n}$ is non-empty for each $n\in\mathbb{N}$, ensuring that we can choose a point $p_{n}\in F_{n}\cap K_{n}^{c}$. To prove that $p_{n}$’s form a discrete subset of $X$ as needed by our strategy, assume to get a contradiction that there exists a subsequence $(p_{n_{k}})$ converging to $p$. This implies that there exists $N\in\mathbb{N}$ such that the $N$-th compact captures our sequence eventually: $p_{n_{k}}\in K_{N}$ for all $n_{k}$; but that contradicts the choice of $p_{n}$, finishing the proof that $p_{n}$ has no limit point in $X$. Thus $C:=\{p_{n}:n\in\mathbb{N}\}$ is a closed subset of $X$ whose image $\rho(C)=\{z_{n}:n\in\mathbb{N}\}$ fails to be closed as its limit point $z_{0}$ lies outside it, establishing that $\rho$ is not a closed map. ∎
We turn to questions which arise when we have more than one retraction map and some initial remarks about the basic question about how do we generate new retraction maps out of old. Algebraic or other operations would require special assumptions about the ambient space; but whatever be the ambient space $X$, viewing retraction maps as self-maps of $X$, we may always compose any two such maps and ask when does it result in a retraction? As the first answer, we have: if two retraction maps $\rho_{1},\rho_{2}$ on $X$ commute, then $\rho:=\rho_{1}\circ\rho_{2}=\rho_{2}\circ\rho_{1}$ is also a retraction map on $X$ whose image is the retract given by the intersection of the retracts $\rho_{1}(X)$ and $\rho_{2}(X)$. If we are allowed to change the space, then there are even more trivial ways such as the following. Given any retraction with notations as above, we may pick an arbitrary subset $A$ of the retract $Z$ and consider the restriction of $\rho$ to $X\setminus\rho^{-1}(A)$; this renders $Z\setminus A$ as a retract of $X\setminus\rho^{-1}(A)$. In other words, we may get retracts of $X$ as intersections of retracts of a space containing $X$ with $X$. The question of whether all retracts of $X\setminus A$ can be realized as being obtained this way is interesting only when $X$ and $A$ satisfy certain conditions i.e., the converse of the last statement indeed fails and there are nice counterexamples towards this. To put down the simplest of examples in the complex analytic category, consider $X=\mathbb{C}^{2}\setminus\{\rm w-axis\}$ where the standard coordinates are herein and henceforth in this article, denoted by $(z,w)$. We then observe that the retract $Z$ described by the equation $zw=1$ which is the image of the retraction $(z,w)\mapsto(z,1/z)$, cannot be realized as mentioned just earlier – the proof of this is recorded in remark 8.2 following a lemma to answer the aforementioned question in the positive direction.
In passing, we put down a rudimentary fact to keep in the background about comparing retracts of a pair of related spaces; we include its proof alongwith statement despite its obviousness (as with certain other facts), owing to its use for the rest of this article.
@@
In passing, we may ofcourse note that by using the well-known theorem of Kaup–Upmeier [Kaup-Upm], we may replace the polydisc in the above by various other product domains which unlike polydiscs are neither convex nor homogeneous. Specifically, we may take a product $B_{1}\times\ldots\times B_{k}$ (with $k\geq 2$) of bounded balanced pseudoconvex domains with the property that one of the factors admits a non-linear holomorphic map into another, to ensure the existence of non-linear retracts, thereby the proof of the above corollary to go through. There are of-course better proofs and our reason for discussing this proof is to simultaneously make a few general remarks about non-linear retracts and about some simple ways of generating ample examples of retracts. We note that there are plenty of product domains of the kind just mentioned (satisfying the property about the existence of non-linear holomorphic maps, as mentioned above). To indicate a way of generating examples, let $E_{2m}:=\{(z,w)\;:\;|z|^{2m}+|w|^{2}<1\}$. For an arbitrary point $(z,w)\in E_{2m}$ note that $|z|<1$, thereby that, $|z^{n}|^{2l}+|w|^{2}<|z|^{2m}+|w|^{2}$ for every $l\geq m$. Consequently, $(z^{n},w)$ lies in $E_{2l}$ for every $l\geq m$, implying that $(z,w)\mapsto(z^{n},w)$, gives an example of a non-linear holomorphic mapping as soon as $n\geq 2$ as desired and more importantly thereby, a non-linear retract passing through the origin in the polyball $E_{2m}\times E_{2l}$ for every $l\geq m$. Needless to say, similar constructions can be done with $l_{p}$-balls and more; we shall not digress further along this line.
@@ 2.2. Notions of extremality and strict convexity
Let $K$ be the field $\mathbb{R}$ or $\mathbb{C}$. Let $A\subset\mathbb{C}^{N}$ be an arbitrary set and $a\in\partial{A}$. We introduce a notion that captures the set of all directions along which one can move from $a$, at-least a little bit while staying within the boundary $\partial A$. Towards this, define the face of $\partial A$ at $a$ by
$F_{K}(a,\partial{A})=\{u\in\mathbb{C}^{N}:\exists\epsilon>0,a+tu\in\partial{A},\text{ for all }t\in K~{}with~{}|t|<\epsilon\}$
reduces to $\{0\}$. It is immediate that $F_{K}$ is closed under scalar multiplication. But a stronger claim that $F_{K}$ is a linear subspace of $\mathbb{C}^{N}$ requires more assumptions on $A$. Indeed, if $A$ is convex set, then $F_{K}(a,\partial{A})$ is $K$-linear subspace. We shall mostly deal with the case when $A=\Omega$ is a domain. In such a case, by abuse of notation we shall write $F(a,\partial\Omega)$ or $F(a,\overline{\Omega})$ in place of $F(a,\Omega)$. A boundary point $a\in\partial{\Omega}$ is said to be a weakly  $K$-extremal boundary point for $\Omega$, if $F_{K}(a,\partial{\Omega})=0$; if the face reduces to $\{0\}$ for every $a\in\partial{\Omega}$, then we say that the domain $\Omega$ is weakly $K$-extremal.
A boundary point $a\in\partial{\Omega}$ is said to be a strongly  $K$-extremal boundary point for $\Omega$, if the face
$F_{K}(a,\overline{\Omega})=\{u\in\mathbb{C}^{N}:\exists\epsilon>0,a+tu\in\overline{\Omega},\text{ for all }t\in K~{}with~{}|t|<\epsilon\}$
reduces to $\{0\}$. As is usual, if all boundary points are strongly $K$-extremal, then we say that the domain itself is strongly $K$-extremal. In general, if $\Omega$ is strongly $K$-extremal, then we shall drop the adjective ‘strongly’ and call $\Omega$ simply as a $K$-extremal domain.
We now prove that in the case of bounded balanced pseudoconvex domains, strongly $\mathbb{C}$-extremal and weakly $\mathbb{C}$-extremal are equivalent, thereby we can forget about the qualifications ‘weak’ and ‘strong’ essentially for the rest of this article. We leave the easier implication about strong extremality implying weak extremality aside and detail only the relatively non-trivial reverse implication.
Proposition 2.13.
Suppose $D$ is a bounded balanced pseudoconvex domain and $p\in\partial D$ is a weakly $\mathbb{C}$-extremal boundary point. Then it is also a strongly $\mathbb{C}$-extremal boundary point.
Proof.
Assume that $p$ is not strongly $\mathbb{C}$-extremal. Then there exists $\epsilon_{0}>0$ and non-zero $u\in\mathbb{C}^{N}$ such that $p+tu\in\overline{D}$ for all $t\in\mathbb{C}$ with $|t|<\epsilon_{0}$. After a rescaling, we may assume that $\epsilon_{0}=1$. Then the affine (holomorphic) map $f$ defined by $f(t)=p+tu$ maps $\Delta$ into $\overline{D}$; note that $f(0)=p$. Let $h$ denote the Minkowski functional of $D$. Consider the map $\varphi:\Delta\to\mathbb{R}$ defined by $\varphi(z)=h\circ f(z)$. Note that $\varphi$ is a plurisubharmonic function whose modulus attains maximum at $z=0$. By maximum principle of plurisubharmonic functions, $\varphi$ is identically constant, which is equal to $1$. This means that $h\circ f\equiv 1$. So ${\rm image}(f)\subset\partial D$. Hence $p$ is not weakly $\mathbb{C}$-extreme boundary point. This finishes the proof by contradiction. ∎
Remark 2.14.
If $K=\mathbb{R}$, then in the case of bounded balanced convex domains, the notion of weak $\mathbb{R}$-extremality and strong $\mathbb{R}$-extremality are equivalent.
@@
$\vec{F}_{\mathbb{C}}(a,\overline{\Omega})=\vec{F}_{\mathbb{R}}(a,\overline{\Omega})\cap i\vec{F}_{\mathbb{R}}(a,\overline{\Omega}).$
It follows that if $\Omega$ is weakly (resp. strongly) $\mathbb{R}$-extremal, then $\Omega$ is weakly (resp. strongly) $\mathbb{C}$-extremal. But the converse is not true. To see this, consider $B_{\ell^{1}}:=\{(z,w)\in\mathbb{C}^{2}:|z|+|w|<1\}$; we shall presently proceed to showing that this is $\mathbb{C}$-extremal but not $\mathbb{R}$-extremal. Perhaps even more basically, we must remark that (vacuously) every domain in $\mathbb{C}$ is $\mathbb{C}$-extremal whereas only the strictly convex ones are $\mathbb{R}$-extremal. To discuss a ‘non-vacuous’ example, we discuss the well-known $\ell^{1}$-ball as mentioned above whose proof is included here owing to its basic importance for us.
@@
We first show that $B_{\ell^{1}}$ is $\mathbb{C}$-extremal. To do this, let $a\in\partial B_{\ell^{1}}$. Suppose there exists $\epsilon_{0}>0$ and $u_{0}\in\mathbb{C}^{N}$ such that $a+tu_{0}\in\partial B_{\ell^{1}}$, for all $t\in\mathbb{C}$ with $|t|<\epsilon_{0}$.
@@
We shall now discuss the relationship between the above extremality notions and that of strict convexity.
@@
In order to establish equivalence of $\mathbb{R}$-extremality with strict convexity, it is convenient to do this, the intermediate concept namely of rotundity, which is defined as follows:
Definition 2.18.
[…]
First we prove that the notion of rotundity and strict convexity are equivalent in the realm of bounded balanced convex domains.
@@
Assume $\Omega$ is rotund. Let $a\in\partial\Omega$ and assume that there exists $u\in\mathbb{C}^{N}$ such that $a+tu\in\partial\Omega$ for all $t\in\mathbb{R}$ with $|t|<\epsilon$. Note that $\frac{1}{2}(a+(a+tu))=a+\frac{t}{2}u\in\partial\Omega$ for all $t\in\mathbb{R}$ with $|t|<\epsilon$. By proposition 5.1.2 of [Meggn], $a=a+tu$ for all $t\in\mathbb{R}$ with $|t|<\epsilon$ and this implies that $u=0$. Therefore, $\Omega$ is $\mathbb{R}$-extremal.
Conversely assume that $\Omega$ is $\mathbb{R}$-extremal. Let $z_{1},z_{2}\in\partial\Omega$.
@@
It follows that when $|t|\leq 1/2$, we have
$\frac{z_{1}+z_{2}}{2}+t(z_{2}-z_{1})=(1/2-t)z_{1}+(1/2+t)z_{2}\in\partial{\Omega}.$
Therefore $z_{2}-z_{1}\in F_{\mathbb{R}}(\frac{z_{1}+z_{2}}{2},\partial{\Omega})$. Since $\Omega$ is $\mathbb{R}$- extremal, $z_{2}=z_{1}$ and therefore $\Omega$ is rotund. ∎
As the relation between the pair of notions gotten by taking $K=\mathbb{C}$ versus $K=\mathbb{R}$ will be basic importance in the sequel, we discuss this below. But one would perhaps like to ascertain first the relation between strict convexity and extremality introduced above. This follows from known facts (see for instance [Meggn]), but owing to its central importance for us, we include the details.
@@
Suppose that $\Omega$ is strictly $\mathbb{R}$-convex. If possible, assume that $\Omega$ is not $\mathbb{R}$- extremal. Then by theorem 2.19, $\Omega$ is not rotund. By corollary 5.1.16 of [Meggn], there exists $\mathbb{R}$-linear functional $f$ supports more than one point. This implies that there exists $z_{0},w_{0}\in\overline{\Omega}$ such that $f(z_{0})=f(w_{0})=\|f\|$. Take $\alpha:=\|f\|$. Consider the $\mathbb{R}$-linear functional $g$ defined by $g(z)=\frac{1}{\alpha}f(z)$. Note that $\|g|\ =1$ and $g(z_{0})=g(w_{0})=1$. This contradicts the strict $\mathbb{R}$-convexity of $\Omega$. Therefore, $\Omega$ is $\mathbb{R}$- extremal.
For the converse, assume that $\Omega$ is $\mathbb{R}$-extremal. Let $T$ be a $\mathbb{R}-$ linear map with $\|T\|=1$. Suppose that there exist $z,w\in\overline{\Omega}$ such that $Tz=1=Tw$. Then $z,w\in\partial{\Omega}$, otherwise $T\left(z/\|z\|\right)=1/\|z\|>1$. Note that $T((1-t)z+tw)=1$. Then $(1-t)z+tw\in\partial{\Omega}$ for all $t\in[0,1]$. In particular, if $t=1/2$, we get $\frac{1}{2}(z+w)\in\partial{\Omega}$. Note then that
@@
Then $w-z\in F_{\mathbb{R}}(\frac{z+w}{2},\partial{\Omega})$. Since $\Omega$ is $\mathbb{R}$-extremal, it follows that $w=z$. Therefore $\Omega$ is strictly $\mathbb{R}$-convex. ∎
@@
Conversely, assume that $\Omega$ is strictly $\mathbb{C}$-convex. We have to prove that $\Omega$ is strictly $\mathbb{R}$-convex. By above theorem, it suffices to show that $\Omega$ is $\mathbb{R}$-extremal. If possible assume that there exists $u\in\mathbb{C}^{N},~{}a\in\partial\Omega$ and an $\epsilon>0$ such that $a+tu\in\partial\Omega$ for all $t\in\mathbb{R}$ with $|t|<\epsilon$. Let $M=span_{\mathbb{C}}\{u\}$. We split the proof now into two steps, formulated as ‘claims’.